# New Sorting Algorithm Wave Sort (W-Sort)

### v2

@@ 3. Analysis
4. Experimental Results
Experimental code was run on randomly shuffled sequences and compared to quicksort. Results demonstrate that the basic wave sort algorithm performs close to the theoretical minimum comparison limit of $log_{2}(n!)$. Across multiple experiments, wave sort required approximately $1\%$ more comparisons than this theoretical limit on average. This efficiency gap decreases as sequence length increases (See table 1).
@@ 4. Experimental Results
$2^{20}$ 19,680,967 26,070,556 19,458,756 19,922,945
$2^{21}$ 41,459,069 55,091,250 41,014,653 41,943,041
$2^{22}$ 87,113,082 115,662,060 86,223,598 88,080,385
$2^{23}$ 182,613,766 243,388,480 180,835,793 184,549,377
Table 1. Average Comparison
Though wave sort requires fewer comparisons than quicksort, it performs more data swaps and requires additional block swaps. In wave sort, each block swap moves each element once and minimizes access to the temporary variable to only once in most cases. Each swap is equivalent to approximately 3 block swaps. The total data movement in basic wave sort is more than twice that of quicksort (See table 2).
@@
Table 2. Average Swap
In practice, many sequences encountered are presorted or partially presorted. Wave sort exhibits its worst-case performance when processing fully presorted sequences (See table 3).
@@
Table 3. Worst Comparison
5. Adaptive Improvement
@@
For sorted or partially sorted sequences, regardless of whether they are reversed, an additional check is incorporated into the up-wave to identify the longest presorted sequence as the sorted portion. For a fully sorted sequence, this results in only $n-1$ comparisons. If the sequence is completely reversed, the number of swaps is reduced to $\frac{n}{2}$. This modification effectively short-circuits the processing of presorted sequences (see Listing  8 and table 4).
1func (s seq) upwave(start, end int) {
2 if start == end {
3 return
4 }
5 sortedStart := s.preSorted(start, end)
6 sortedLength := end - sortedStart + 1
7 leftBound := end - sortedLength<<1 + 1
8 length := end - start + 1
9 for leftBound > start {
10 s.downwave(leftBound, sortedStart, end)
11 sortedStart = leftBound
12 sortedLength = end - sortedStart + 1
13 if length < sortedLength<<2 {
14 break
15 }
16 leftBound = end - sortedLength<<1 + 1
17 }
18 s.downwave(start, sortedStart, end)
19}
20
21func (s seq) preSorted(start, end int) int {
22 du := false
23 i := end
24 for i > start {
25 if du {
26 if s.less(i, i-1) {
27 return I
28 }
29 i--
30 continue
31 }
32 if s.less(i-1, i) {
33 if i == end {
34 du = true
35 i--
36 continue
37 }
38 s.reverse(i, end)
39 return i
40 }
41 i--
42 }
43 if du {
44 return start
45 }
46 s.reverse(start, end)
47 return start
48}
49
50func (s seq) reverse(start, end int) {
51 i := start
52 j := end
53 for range (end - start + 1) / 2 {
54 s.swap(i, j)
55 i++
56 j--
57 }
58}
Listing 8: Adaptive Wave Sort
@@
Table 4. Adaptive Presorted
As the adaptive approach adds one more step to check the boundary of the sorted sequence, the number of comparisons increases slightly. This approach trades off swaps for block swaps. The comparisons increase by approximately $0.6\%$. The number of swaps decreases by approximately $1\%$, while block swaps decrease by approximately $7\%$ (See table 5).
@@
$2^{20}$ 19,824,134 4,972,230 15,569,767
$2^{21}$ 41,729,429 10,458,356 33,523,711
$2^{22}$ 87,669,368 21,975,151 70,442,223
$2^{23}$ 183,615,126 45,971,847 152,140,092
Table 5. Adaptive Average
6. Trade off Comparison for Swap
The average comparison count of wave sort is quite close to the theoretical limit, leaving only a small margin for improvement. In practice, comparisons usually cost more than swaps on complex data structures, as comparisons require examining the content of the data structure while swaps only require exchanging pointers to the content. Therefore, fewer comparisons are generally preferred. However, in some cases, trading comparisons for swaps may be advantageous.
Since wave sort performs more than twice the swaps compared to quicksort, there is room for optimization if such trade-offs prove beneficial. The basic strategy is to increase the expansion rate in the up-wave so that partitioning works more effectively to reduce swaps. In both down-wave and up-wave processes, sorting short sequences with insertion sort can replace swaps with block swaps for greater efficiency. The smaller sorted portions also reduce the number of block swaps required. An implementation showing these trade-offs is present in Listing  9
1func (s seq) upwave(start, end int) {
2 length := end - start + 1
3 if length == 1 {
4 return
5 }
6 if length <= 8 {
7 s.insertSort(start, end)
8 return
9 }
10 sortedStart := s.preSorted(start, end)
11 sortedLength := end - sortedStart + 1
12 sortedLength <<= 4
13 leftBound := end - sortedLength + 1
14 for leftBound > start {
15 s.downwave(leftBound, sortedStart, end)
16 sortedStart = leftBound
17 sortedLength = end - sortedStart + 1
18 if length <= sortedLength<<6 {
19 break
20 }
21 sortedLength <<= 4
22 leftBound = end - sortedLength + 1
23 }
24 s.downwave(start, sortedStart, end)
25}
26
27func (s seq) downwave(start, sortedStart, end int) {
28 unsortedLength := sortedStart - start
29 if unsortedLength == 0 {
30 return
31 }
32 lr := end - sortedStart
33 if lr == 0 {
34 if unsortedLength < 8 {
35 s.insertSort(start, end)
36 return
37 }
38 s.upwave(start, end)
39 return
40 }
41 p := sortedStart + lr/2
42 ...
43}
44
45func (s seq) insertSort(l, r int) {
46 sortedIndex := l + 1
47 for sortedIndex <= r {
48 i := sortedIndex
49 for i > l {
50 if s.less(sortedIndex, i-1) {
51 i--
52 continue
53 }
54 break
55 }
56 if i < sortedIndex {
57 tmp := s[sortedIndex]
58 for j := sortedIndex; j > i; j-- {
59 s[j] = s[j-1]
60 }
61 s[i] = tmp
62 }
63 sortedIndex++
64 }
65}
Listing 9: Trade-off Comparison
A typical trade-off result is shown in Table 6 with the implementation demonstrated in the code above. One comparison trades for approximately 4 swaps plus 80 block swaps, which is equivalent to approximately 31 swaps (See table 6). More detailed mathematical analysis is required to determine the optimal expansion rate. Generally, the longer the sorted portion, the larger the expansion rate should be, based on the probability equation provided above.
sequence size, $n$ comparison swap block swap
$2^{20}$ 19,853,004 4,467,539 4,459,874
$2^{21}$ 41,996,088 9,462,760 7,691,231
$2^{22}$ 88,300,820 19,888,658 14,535,769
$2^{23}$ 185,437,750 41,802,470 26,231,740
Table 6. Trade off Comparison for Swap
7. Conclusion
In this paper, we introduced Wave Sort, a novel divide-and-conquer in-place sorting algorithm designed to address key limitations found in widely used comparison sorts. By employing a dynamic pivot selection strategy that leverages an expanding sorted region within the array, Wave Sort overcomes the pivot inconsistency issues that can lead to $O(N^{2})$ worst-case performance in quicksort. Unlike standard merge sort, Wave Sort achieves sorting in-place, avoiding the need for $O(N)$ auxiliary space, and ensures a consistent, low logarithmic recursion stack depth without the need for explicit monitoring, presenting an advantage over algorithms like Introsort in memory-constrained environments.
Our analysis indicates a worst-case comparison complexity of $O(N(\log N)^{2})$ with a small constant factor ($\approx\frac{1}{4}$). Experimental evaluation on randomly shuffled data demonstrated that Wave Sort achieves an average comparison count significantly lower than quicksort (approximately 24% less) and remarkably close to the theoretical minimum of $O(N\log N)$, aligning with our analytical expectations. Furthermore, the algorithm is readily adaptable to handle partially or fully presorted and reversed sequences with high efficiency, achieving linear time complexity in these best-case scenarios through the integrated adaptive improvement. While the basic implementation involves more data movement operations than quicksort, we showed that the algorithm’s structure allows for effective trade-offs to reduce swaps by increasing comparison operations, depending on the relative costs in a given application context.
Wave Sort presents a compelling alternative for scenarios requiring efficient in-place sorting with predictable performance characteristics. Its combination of guaranteed logarithmic stack depth, favorable average comparison performance, and bounded worst-case behavior positions it as a strong candidate for replacing algorithms like in-place merge sort variants and Introsort in standard library implementations, especially where robust and consistent sorting is paramount.
Future work includes conducting a more detailed formal analysis to determine the optimal parameters for trading comparisons against element moves across various data types and hardware architectures, thereby maximizing practical performance. Further research could also explore variations in the initial sorted segment size, alternative median or pivot selection strategies from the sorted portion, potential parallelization of the algorithm, and its applicability to related sorting problems or data structures. Wave Sort represents a promising direction for advancing the state of the art in in-place comparison sorting by effectively balancing theoretical elegance with practical performance considerations.
Acknowledgements.
For my father, Wei Wei, who motivated me to write this, and Professor Ghosal for his very helpful stats class.

### v3

@@ 2 Analysis
3 Experimental Results
The experimental code was run on randomly shuffled sequences and compared to quick sort. The sequence is shuffled with the Shuffle function from the go package ”math/rand/v2”. The quick sort is the improved version with the partition function proposed in this paper which reduces comparisons and keeps the same swaps. The results demonstrate that the basic wave sort algorithm performs close to the theoretical minimum comparison limit of $\log_{2}(n!)$. Across multiple experiments, wave sorting required approximately $1\%$ more comparisons than this theoretical limit on average. This efficiency gap decreases as the sequence length increases (see Table 1). The code is shown in the Appendix A.
@@ 3 Experimental Results
$2^{20}$ 19,680,967 26,070,556 19,458,756 19,660,801
$2^{21}$ 41,459,069 55,091,250 41,014,653 41,418,753
$2^{22}$ 87,113,082 115,662,060 86,223,598 87,031,809
$2^{23}$ 182,613,766 243,388,480 180,835,793 182,452,225
Table 1: Average Comparison
Although wave sort requires fewer comparisons than quick sort, it performs more data swaps and requires additional block swaps. In wave sort, each block swap moves each element once and minimizes access to the temporary variable to only once in most cases. Each swap is equivalent to approximately 3 block swaps. The total data movement in the basic wave sort is more than twice that in the quick sort (see Table 2). This issue is addressed in trade-off comparisons with swaps.
@@
Table 2: Average Swap
In practice, many sequences encountered are presorted or partially presorted. Wave sort exhibits its worst-case performance when processing fully presorted sequences (see Table 3). This issue is addressed in adaptive and hybrid sorting.
@@
Table 3: Worst Comparison
4 Trade off Comparisons with Swaps
The average comparison count of wave sort is quite close to the theoretical limit, leaving only a small margin for improvement. In practice, comparisons usually cost more than swaps on complex data structures, as comparisons require examining the content of the data structure, while swaps only require exchanging pointers to the content. Therefore, in general, fewer comparisons are preferred. However, in some cases, trading comparisons for swaps may be advantageous.
Since wave sort performs more than twice the swaps compared to quick sort, there is room for optimization if such trade-offs prove beneficial. The basic strategy is to increase the expansion rate in the up-wave so that partitioning works more effectively to reduce swaps. In both down and up-wave processes, sorting short sequences with insertion sort can replace swaps with block swaps for greater efficiency. The smaller sorted portions also reduce the number of block swaps required. An implementation showing these trade-offs is present in Listing 8.
func (s seq) upwave(start, end int) {
length := end - start + 1
if length == 1 {
return
if length <= 12 {
s.insertSort(start, end)
return
sortedStart := end - 7
s.insertSort(sortedStart, end)
sortedLength := 8
leftBound := end - 127
for leftBound > start {
s.downwave(leftBound, sortedStart, end)
sortedStart = leftBound
sortedLength = end - sortedStart + 1
if length <= sortedLength<<6 {
break
sortedLength <<= 4
leftBound = end - sortedLength + 1
s.downwave(start, sortedStart, end)
func (s seq) downwave(start, sortedStart, end int) {
unsortedLength := sortedStart - start
if unsortedLength == 0 {
return
if end - sortedStart == 0 {
if unsortedLength < 12 {
s.insertSort(start, end)
return
s.upwave(start, end)
return
p := sortedStart + (end-sortedStart)/2
func (s seq) insertSort(l, r int) {
sortedIndex := r - 1
for ; sortedIndex >= l; sortedIndex-- {
low := sortedIndex
hi := r
m := (low + hi + 1) >> 1
for low < hi {
if s.less(sortedIndex, m) {
hi = m - 1
m = (low + hi + 1) >> 1
continue
low = m
m = (low + hi + 1) >> 1
if m > sortedIndex {
tmp := s[sortedIndex]
for j := sortedIndex; j < m; j++ {
s[j] = s[j+1]
s[m] = tmp
bs += m - sortedIndex + 1
Listing 8: Trade-off Comparison
A typical trade-off result is shown in Table 4 with the implementation demonstrated in the above code. The comparison increases approximately $1.6\%$. One comparison trades for approximately 2.6 swaps plus 46 block swaps, which is equivalent to approximately 25 swaps (see Table 4). The code is shown in the Appendix B. More detailed mathematical analysis is required to determine the optimal expansion rate. Generally, the longer the sorted portion, the larger the expansion rate should be, based on the probability equation provided above.
sequence size, $n$ comparison swap block swap
$2^{20}$ 20,105,673 4,235,207 4,252,992
$2^{21}$ 42,359,965 8,974,366 9,777,278
$2^{22}$ 87,822,911 19,070,908 22,700,097
$2^{23}$ 185,053,654 40,244,980 36,335,453
Table 4: Trade off Comparison for Swap
5 Adaptive and Hybrid Sorting
@@
For sorted or partially sorted sequences, regardless of whether they are reversed, an additional check is incorporated into the up-wave to identify the longest presorted sequence as the sorted portion. For a fully sorted sequence, this results in only n1n-1 comparisons. If the sequence is completely reversed, the number of swaps is reduced to $\frac{n}{2}$. This modification effectively short-circuits the processing of presorted sequences (see Listing 9 and Table 5). The code is shown in the Appendix C.
func (s seq) upwave(start, end int) {
length := end - start + 1
if length == 1 {
return
if length <= 12 {
s.insertSort(start, end, end)
return
sortedStart := s.preSorted(start, end)
sortedLength := end - sortedStart + 1
if sortedLength < 8 {
s.insertSort(end-7, sortedStart, end)
sortedStart = end - 7
sortedLength = 8
sortedLength <<= 4
leftBound := end - sortedLength + 1
for leftBound > start {
s.downwave(leftBound, sortedStart, end)
sortedStart = leftBound
sortedLength = end - sortedStart + 1
if length <= sortedLength<<6 {
break
sortedLength <<= 4
leftBound = end - sortedLength + 1
s.downwave(start, sortedStart, end)
func (s seq) preSorted(start, end int) int {
du := false
for i := end; i > start; i-- {
if du {
if s.less(i, i-1) {
return i
continue
if s.less(i-1, i) {
if i == end {
du = true
continue
s.reverse(i, end)
return i
if du {
return start
s.reverse(start, end)
return start
func (s seq) reverse(start, end int) {
i := start
j := end
for range (end - start + 1) / 2 {
s.swap(i, j)
i++
j--
func (s seq) insertSort(l, sortedStart, r int) {
sortedIndex := sortedStart - 1
for ; sortedIndex >= l; sortedIndex-- {
i := sortedIndex
for ; i < r; i++ {
if s.less(sortedIndex, i+1) {
break
if i > sortedIndex {
tmp := s[sortedIndex]
for j := sortedIndex; j < i; j++ {
s[j] = s[j+1]
s[i] = tmp
Listing 9: Adaptive Wave Sort
@@
Table 5: Adaptive Presorted
As the adaptive approach adds one more step to check the boundary of the sorted sequence, the number of comparisons increases slightly. This approach trades off swaps for block swaps. The comparisons increase by approximately $0.15\%$. The number of swaps increases by approximately $0.53\%$, while the block swaps decrease by approximately $1.6\%$ (see Table 6).
@@
$2^{20}$ 20,140,905 4,261,710 4,170,009
$2^{21}$ 42,416,613 9,017,588 9,642,118
$2^{22}$ 87,949,166 19,166,119 22,400,068
$2^{23}$ 185,336,094 40,456,208 35,665,233
Table 6: Adaptive Average
The up-wave could adopt other adaptive sorting to short-circuit certain sequence patterns, such as shown in MBISort MBISort. The best adaptive sorting needs further study to ensure that it does not introduce too many extra comparisons or swaps in not-covered patterns.
As adaptive sorting cannot cover all the possible sequence patterns, there are still sequence patterns for the wave sort to hit the worst case. The heap sort could be taken as the last resort, as in Introsort introsort and pdqsort pdqsort during the down-wave operation to ensure $O(N\log(N))$ comparisons in the worst case. From the distribution of pivot position, as shown in figure 5 (it demonstrates the probability of pivot position in the case of 99 elements plus the pivot element. The wave sort has 99 samples and the median of three has 3 samples in this case), the probability of pivot position in wave sorting in the range of 1 to $\frac{1}{8}n$ is very low, while the probability of median of three in the same range as in pdqsort pdqsort is much higher. So in wave sort, it is quite unlikely that it falsely falls in the heap sort too early. It is also safe for Introsort to take more aggressive approaches to switch to heap sort earlier if there are unbalanced partitions occurring.
Figure 5: pivot-distribution
6 Conclusion
In this paper, we introduced Wave Sort, a novel divide-and-conquer in-place sorting algorithm designed to address key limitations found in widely used comparison sorts. Using a dynamic pivot selection strategy that leverages an expanding sorted region within the array, Wave Sort overcomes the pivot inconsistency issues that can lead to $O(N^{2})$ worst-case performance in quicksort. Unlike standard merge sort, Wave Sort achieves sorting in place, avoiding the need for $O(N)$ auxiliary space, and ensures a consistent, low logarithmic recursion stack depth. We refined the partition algorithm used by many standard quicksort libraries to reduce comparisons from $N+1$ to N1N-1, the theoretical limit, on an array of size $N$. The performance of quick sorting is improved. We proposed a novel block-swap algorithm to reduce the movement of elements to the theoretical limit. It could be modified to replace the rotating swap mentioned in in-place-mergesort. The block swap performs $\frac{a+b}{2}$ swaps on arrays of sizes $a$ and $b$. The rotating swap may perform $a+b$ swaps on the same arrays. With all the extra operations removed, it helps to improve current standard libraries and performance analysis.
The algorithm is divided into two logic components, the up- and down-wave, each of which could have different expansion approaches, such as integration with adaptive sorting or hybrid with other sorting algorithms.
Our analysis indicates a worst-case comparison complexity of $O(N(\log(N))^{2})$ with a small constant factor ($\approx\frac{1}{4}$). Experimental evaluation of randomly shuffled data demonstrated that Wave Sort achieves an average comparison count significantly lower than quicksort (approximately $24\%$ less) and remarkably close to the theoretical minimum of $\log(N)!)$, aligning with our analytical expectations. Furthermore, the algorithm is readily adaptable to handle partially or fully presorted and reversed sequences with high efficiency, achieving linear-time complexity in these best-case scenarios through integrated adaptive improvement. While the basic implementation involves more data movement operations than quicksort, we showed that the algorithm’s structure allows for effective trade-offs to reduce swaps by increasing comparison operations, depending on the relative costs in a given application context. Running on a laptop with the comparison of integers, which is the fastest data type to compare in most CPU architectures and is implemented in the Go programming language, sorting a one million integer sequence takes about $12.96$ seconds for 100 runs. The improved quicksort takes about $11.71$ seconds. After increasing the complexity of the comparison, the wave sort takes about $120$ seconds and the improved quicksort takes about $140$ seconds. As the code is mainly for showcasing the algorithm and is not optimized for performance, and since Go is not an ideal language for heavy lifting jobs, the running time is just for reference.
Wave Sort presents a compelling alternative for scenarios that require efficient in-place sorting with predictable performance characteristics. Its combination of guaranteed logarithmic stack depth, favorable average comparison performance, and bounded worst-case behavior positions it as a strong candidate for replacing algorithms like in-place merge sort variants and the quick sort in Introsort in standard library implementations, especially where robust and consistent sorting is paramount.
Future work includes conducting a more detailed formal analysis to determine the optimal parameters for trading comparisons against element moves, thereby maximizing practical performance. Further research could also explore variations in adaptive sorting, hybrid sorting in Introsort, and its applicability to related sorting problems or data structures. Wave Sort represents a promising direction for advancing the state-of-the-art in in-place comparison sorting by effectively balancing theoretical elegance with practical performance considerations.
Acknowledgments
To my father, Wei Wei, for inspiring me to write this in the first place.
Appendix A wave sort - basic
package wave_sort
type seq []int
var compare, swap, blockSwap int
// partition splits the sequence into two parts, the left part is less than the pivot and the right part is greater than the pivot.
// It returns the index of the split point which is involved in the greater part.
// l is the index of the left bound of the unsorted part.
// r is the index of the right bound of the unsorted part which is the firt element of the sorted part.
// p is the index of the pivot.
func (s seq) partition(l, r, p int) int {
i, j := l-1, r
for {
for {
i++
if i == j {
return i
if s.less(p, i) {
break
for {
j--
if j == i {
return i
if s.less(j, p) {
break
s.swap(i, j)
func (s seq) less(i, j int) bool {
compare++
return s[i] < s[j]
// blockSwap swaps the block from m to r and the one from r to p.
// m is the index of the first element of the left block.
// r is the index of the right bound of the left block and the first element of the right block.
// p is the index of the last element of the right block.
func (s seq) blockSwap(m, r, p int) {
$l$ := r - m
if $l$ == 0 {
return
lr := p - r + 1
if lr == 1 {
s[m], s[p] = s[p], s[m]
swap++
return
if lr <= $l$ {
s.blockSwap_sr(m, r, p)
blockSwap += lr << 1
return
s.blockSwap_sl(m, p, $l$)
blockSwap += $l$ + lr
func (s seq) blockSwap_sl(m, p, $l$ int) {
tmp := s[m]
init := m
j := m
nm := p - $l$ + 1
var k int
for range p - m + 1 {
if j >= nm {
k = j - nm + m
if k == init {
init++
s[j] = tmp
j = init
tmp = s[j]
continue
} else {
k = j + $l$
s[j] = s[k]
j = k
func (s seq) blockSwap_sr(m, r, p int) {
i := m
tmp := s[i]
j := r
for j < p {
s[i] = s[j]
i++
s[j] = s[i]
j++
s[i] = s[j]
s[j] = tmp
func (s seq) swap(i, j int) {
swap++
s[i], s[j] = s[j], s[i]
func (s seq) downwave(start, sortedStart, end int) {
if sortedStart-start == 0 {
return
p := sortedStart + (end-sortedStart)/2
m := s.partition(start, sortedStart, p)
if m == sortedStart {
if p == sortedStart {
s.upwave(start, sortedStart-1)
return
s.downwave(start, sortedStart, p-1)
return
s.blockSwap(m, sortedStart, p)
if m == start {
if p == sortedStart {
s.upwave(m+1, end)
return
p++
s.downwave(m+p-sortedStart, p, end)
return
if p == sortedStart {
s.upwave(start, m-1)
s.upwave(m+1, end)
return
s.downwave(start, m, m+p-sortedStart-1)
s.downwave(m+p-sortedStart+1, p+1, end)
func (s seq) upwave(start, end int) {
if start == end {
return
sortedStart := end
sortedLength := 1
leftBound := end - 1
length := end - start + 1
for leftBound > start {
s.downwave(leftBound, sortedStart, end)
sortedStart = leftBound
sortedLength = end - sortedStart + 1
if length < sortedLength<<2 {
break
leftBound = end - sortedLength<<1 + 1
s.downwave(start, sortedStart, end)
func (s seq) WaveSort() {
if len(s) < 2 {
return
s.upwave(0, len(s)-1)
Listing 10: Wave Sort Basic Implementation
Appendix B wave sort - trade off
package trade_off
type seq []int
var compare, swap, blockSwap int
// partition splits the sequence into two parts, the left part is less than the pivot and the right part is greater than the pivot.
// It returns the index of the split point which is involved in the greater part.
// l is the index of the left bound of the unsorted part.
// r is the index of the right bound of the unsorted part which is the firt element of the sorted part.
// p is the index of the pivot.
func (s seq) partition(l, r, p int) int {
i, j := l-1, r
for {
for {
i++
if i == j {
return i
if s.less(p, i) {
break
for {
j--
if j == i {
return i
if s.less(j, p) {
break
s.swap(i, j)
func (s seq) less(i, j int) bool {
compare++
return s[i] < s[j]
// blockSwap swaps the block from m to r and the one from r to p.
// m is the index of the first element of the left block.
// r is the index of the right bound of the left block and the first element of the right block.
// p is the index of the last element of the right block.
func (s seq) blockSwap(m, r, p int) {
$l$ := r - m
if $l$ == 0 {
return
lr := p - r + 1
if lr == 1 {
s[m], s[p] = s[p], s[m]
swap++
return
if lr <= $l$ {
s.blockSwap_sr(m, r, p)
blockSwap += lr << 1
return
s.blockSwap_sl(m, p, $l$)
blockSwap += $l$ + lr
func (s seq) blockSwap_sl(m, p, $l$ int) {
tmp := s[m]
init := m
j := m
nm := p - $l$ + 1
var k int
for range p - m + 1 {
if j >= nm {
k = j - nm + m
if k == init {
init++
s[j] = tmp
j = init
tmp = s[j]
continue
} else {
k = j + $l$
s[j] = s[k]
j = k
func (s seq) blockSwap_sr(m, r, p int) {
i := m
tmp := s[i]
j := r
for j < p {
s[i] = s[j]
i++
s[j] = s[i]
j++
s[i] = s[j]
s[j] = tmp
func (s seq) swap(i, j int) {
swap++
s[i], s[j] = s[j], s[i]
func (s seq) upwave(start, end int) {
length := end - start + 1
if length == 1 {
return
if length <= 12 {
s.insertSort(start, end)
return
sortedStart := end - 7
s.insertSort(sortedStart, end)
sortedLength := 8
leftBound := end - 127
for leftBound > start {
s.downwave(leftBound, sortedStart, end)
sortedStart = leftBound
sortedLength = end - sortedStart + 1
if length <= sortedLength<<6 {
break
sortedLength <<= 4
leftBound = end - sortedLength + 1
s.downwave(start, sortedStart, end)
func (s seq) downwave(start, sortedStart, end int) {
unsortedLength := sortedStart - start
if unsortedLength == 0 {
return
if end-sortedStart == 0 {
if unsortedLength < 12 {
s.insertSort(start, end)
return
s.upwave(start, end)
return
p := sortedStart + (end-sortedStart)/2
m := s.partition(start, sortedStart, p)
if m == sortedStart {
if p == sortedStart {
s.upwave(start, sortedStart-1)
return
s.downwave(start, sortedStart, p-1)
return
s.blockSwap(m, sortedStart, p)
if m == start {
if p == sortedStart {
s.upwave(m+1, end)
return
s.downwave(m+p-sortedStart+1, p+1, end)
return
if p == sortedStart {
s.upwave(start, m-1)
s.upwave(m+1, end)
return
s.downwave(start, m, m+p-sortedStart-1)
s.downwave(m+p-sortedStart+1, p+1, end)
func (s seq) WaveSort() {
if len(s) < 2 {
return
s.upwave(0, len(s)-1)
func (s seq) insertSort(l, r int) {
sortedIndex := r - 1
for ; sortedIndex >= l; sortedIndex-- {
low := sortedIndex
hi := r
m := (low + hi + 1) >> 1
for low < hi {
if s.less(sortedIndex, m) {
hi = m - 1
m = (low + hi + 1) >> 1
continue
low = m
m = (low + hi + 1) >> 1
if m > sortedIndex {
tmp := s[sortedIndex]
for j := sortedIndex; j < m; j++ {
s[j] = s[j+1]
s[m] = tmp
blockSwap += m - sortedIndex + 1
Listing 11: Wave Sort with Trade-offs
Appendix C wave sort - adaptive
package adaptive
type seq []int
var compare, swap, blockSwap int
// partition splits the sequence into two parts, the left part is less than the pivot and the right part is greater than the pivot.
// It returns the index of the split point which is involved in the greater part.
// l is the index of the left bound of the unsorted part.
// r is the index of the right bound of the unsorted part which is the firt element of the sorted part.
// p is the index of the pivot.
func (s seq) partition(l, r, p int) int {
i, j := l-1, r
for {
for {
i++
if i == j {
return i
if s.less(p, i) {
break
for {
j--
if j == i {
return i
if s.less(j, p) {
break
s.swap(i, j)
func (s seq) less(i, j int) bool {
compare++
return s[i] < s[j]
// blockSwap swaps the block from m to r and the one from r to p.
// m is the index of the first element of the left block.
// r is the index of the right bound of the left block and the first element of the right block.
// p is the index of the last element of the right block.
func (s seq) blockSwap(m, r, p int) {
$l$ := r - m
if $l$ == 0 {
return
lr := p - r + 1
if lr == 1 {
s[m], s[p] = s[p], s[m]
swap++
return
if lr <= $l$ {
s.blockSwap_sr(m, r, p)
blockSwap += lr << 1
return
s.blockSwap_sl(m, p, $l$)
blockSwap += $l$ + lr
func (s seq) blockSwap_sl(m, p, $l$ int) {
tmp := s[m]
init := m
j := m
nm := p - $l$ + 1
var k int
for range p - m + 1 {
if j >= nm {
k = j - nm + m
if k == init {
init++
s[j] = tmp
j = init
tmp = s[j]
continue
} else {
k = j + $l$
s[j] = s[k]
j = k
func (s seq) blockSwap_sr(m, r, p int) {
i := m
tmp := s[i]
j := r
for j < p {
s[i] = s[j]
i++
s[j] = s[i]
j++
s[i] = s[j]
s[j] = tmp
func (s seq) swap(i, j int) {
swap++
s[i], s[j] = s[j], s[i]
func (s seq) upwave(start, end int) {
length := end - start + 1
if length == 1 {
return
if length <= 12 {
s.insertSort(start, end, end)
return
sortedStart := s.preSorted(start, end)
sortedLength := end - sortedStart + 1
if sortedLength < 8 {
s.insertSort(end-7, sortedStart, end)
sortedStart = end - 7
sortedLength = 8
sortedLength <<= 4
leftBound := end - sortedLength + 1
for leftBound > start {
s.downwave(leftBound, sortedStart, end)
sortedStart = leftBound
sortedLength = end - sortedStart + 1
if length <= sortedLength<<6 {
break
sortedLength <<= 4
leftBound = end - sortedLength + 1
s.downwave(start, sortedStart, end)
func (s seq) downwave(start, sortedStart, end int) {
unsortedLength := sortedStart - start
if unsortedLength == 0 {
return
lr := end - sortedStart
if lr == 0 {
if unsortedLength < 12 {
s.insertSort(start, end, end)
return
s.upwave(start, end)
return
p := sortedStart + lr/2
m := s.partition(start, sortedStart, p)
if m == sortedStart {
if p == sortedStart {
s.upwave(start, sortedStart-1)
return
s.downwave(start, sortedStart, p-1)
return
s.blockSwap(m, sortedStart, p)
if m == start {
if p == sortedStart {
s.upwave(m+1, end)
return
s.downwave(m+p-sortedStart+1, p+1, end)
return
if p == sortedStart {
s.upwave(start, m-1)
s.upwave(m+1, end)
return
s.downwave(start, m, m+p-sortedStart-1)
s.downwave(m+p-sortedStart+1, p+1, end)
func (s seq) WaveSort() {
if len(s) < 2 {
return
s.upwave(0, len(s)-1)
func (s seq) preSorted(start, end int) int {
du := false
for i := end; i > start; i-- {
if du {
if s.less(i, i-1) {
return i
continue
if s.less(i-1, i) {
if i == end {
du = true
continue
s.reverse(i, end)
return i
if du {
return start
s.reverse(start, end)
return start
func (s seq) reverse(start, end int) {
i := start
j := end
for range (end - start + 1) / 2 {
s.swap(i, j)
i++
j--
func (s seq) insertSort(l, sortedStart, r int) {
sortedIndex := sortedStart - 1
for ; sortedIndex >= l; sortedIndex-- {
low := sortedIndex
hi := r
m := (low + hi + 1) >> 1
for low < hi {
if s.less(sortedIndex, m) {
hi = m - 1
m = (low + hi + 1) >> 1
continue
low = m
m = (low + hi + 1) >> 1
if m > sortedIndex {
tmp := s[sortedIndex]
for j := sortedIndex; j < m; j++ {
s[j] = s[j+1]
s[m] = tmp
blockSwap += m - sortedIndex + 1
Listing 12: Adaptive Wave Sort